### v1

# Large-scale Generative AI Models Lack Visual Number Sense

Abstract
Humans can readily judge the number of objects in a visual scene, even without counting, and such a skill has been documented in a variety of animal species and in babies prior to language development and formal schooling. Numerical judgments are error-free for small sets, while for larger collections responses become approximate, with variability increasing proportionally to the target number. This response pattern is observed for items of all kinds, despite variation in object features (such as color or shape), suggesting that our visual number sense relies on abstract representations of numerosity. Here, we investigated whether generative Artificial Intelligence (AI) models based on large-scale transformer architectures can reliably name the number of objects in simple visual stimuli or generate images containing a target number of items in the 1-10 range. Surprisingly, none of the foundation models considered performed in a human-like way: They all made striking errors even with small numbers, the response variability often did not increase in a systematic way, and the pattern of errors varied with object category. Our findings demonstrate that advanced AI systems still lack a basic ability that supports an intuitive understanding of numbers, which in humans is foundational for numeracy and mathematical development.
Keywords Foundation Models  $\cdot$ Machine Vision  $\cdot$ Deep Learning  $\cdot$ Numerical Cognition  $\cdot$ Numerosity
Introduction
Artificial Intelligence (AI) is progressing rapidly, with deep learning models approaching or even surpassing human performance in a variety of domains, including perceptual judgements [1] and natural language processing [2]. In this article, we investigate whether advanced AI systems can judge the numerosity of visual sets, a core capability that we share with many animal species [3]. Even infants are sensitive to numerosity [4] and toddlers can generate sets containing a target number of items [5], suggesting that a preverbal understanding of numerical quantities develops well before formal education. Although it is a high-level feature derived from global image properties, numerosity is spontaneously extracted by our visual system [6]. Small numerosities in the “subitizing” range (up to 4) are perceived in an exact manner, while the numerosity of larger sets is approximately estimated when counting is precluded [7]. In the latter case, responses follow Weber’s law, so that variability increases proportionally to the mean estimate [3]. Another key signature of number sense is its abstract nature: numerosity is encoded independently from object category, location or presentation modality [4].
Here we systematically probed the visual number sense of several state-of-the-art “foundation models”, which are large-scale generative architectures trained on huge data sets that can be readily used to solve a wide range of downstream tasks [8]. In the image-to-text domain, modern Visual Question Answering (VQA) systems can provide written answers to non-trivial questions about the content of an image or accurate descriptions of complex visual scenes; here we tested the capability of judging visual numerosity of the Vision-and-Language Transformer (ViLT) [9] and the Bootstrapping Language-Image Pre-training (BLIP-2) model [10]. In the text-to-image domain, generative models can produce high-quality visual content following detailed user prompts provided in natural language; here we tested the numerosity production skills of two of the most popular generative models for images: DALL-E (version 2) [11] and Stable Diffusion (version 2.1) [12].
In line with recent proposals [13], we exploited two behavioral tasks used in cognitive psychology to evaluate number sense: numerosity naming [7], which requires establishing how many items are present in a given image, and numerosity production [14, 5], which requires generating a target number of items. We characterized the distribution of responses produced by the different AI models in the range 1-10 using a variety of object categories. Perfect accuracy across the entire numerical range would suggest the emergence of counting skills, while error-free responses with only small numbers would either indicate subitizing capabilities, or that counting is only partially developed as in children who do not fully master the counting principles [15, 16]. Error-prone responses centered on the target number would instead suggest that the AI model relies on approximate estimation mechanisms, which may or may not follow Weber’s law.
Results
Numerosity naming
We probed the numerosity naming skills of two state-of-the-art VQA architectures: the Vision-and-Language Transformer (ViLT) [9] and the Bootstrapping Language-Image Pre-training (BLIP-2) model [10]. These systems have remarkable visual reasoning abilities and can answer nontrivial questions related to image content (e.g., What does the image represent? What are the feelings of the people in the scene and why?). We asked the models how many objects were present in a set of images containing up to 10 objects (see Section Materials and Methods for methodological details). Each stimulus included only items of the same kind, sampled from 5 possible classes: realistic pictures of common categories (apples, people, butterflies), colored dots, and “fast cards” depicting regularly placed clip-arts similar to those used to test children [15]. Examples are shown in Fig. 1 (top row).
Figure 1: Examples of images provided as input or produced by the models, ordered according to target number (reported at the bottom). The first row shows examples of stimuli used in the numerosity naming task: the first four images are taken from the “apples”, “people”, “butterflies” and “dots” categories, while the remaining images are sampled from the “fast cards” category. Subsequent rows show examples of images generated by DALL-E and Stable Diffusion in the numerosity production task, showcasing both correct and wrong generations. We sampled two images for each target category (apples, people, butterflies, dots). In specific cases, DALL-E never makes mistakes when there are a few objects to be produced (up to 3 objects for the “apples” category, exactly 3 objects for the “people” category, and exactly 1 object for the “butterflies” category). Another interesting phenomenon is the generation of images containing a wrong number of dots, which are nevertheless arranged according to the target digit shape (e.g., 8 in the figure).
In general, response accuracy was extremely low for both models (ViLT: 28.0%, BLIP-2: 29.6%), thereby showing that they cannot count. In sharp contrast with humans, both models returned wrong answers within the subitizing range (1-4) and even for images with one or two objects. According to standard criteria used in developmental studies (see Methods), these models could be considered at most “One”-knowers [15]. Confusion matrices (CMs) reported in Fig. 2 clearly show the presence of anchoring effects, leading the models to choose prototypical responses (e.g., 4, 6) probably corresponding to common spatial layouts that might be over-represented in the training corpora. The pattern of responses also drastically varied between categories (minimum correlation between CMs for ViLT: 0.04, BLIP-2: 0.27), suggesting that the emergent numerical representations are not abstract.
Numerosity production
We also probed the numerosity production skills of two of the most popular generative models for images: DALL-E (version 2) [11] and Stable Diffusion (version 2.1) [12]. These foundation models have proven capable of generating high-quality images following a textual description, also taking into account stylistic instructions, fine-grained details, and relational features (e.g., A photo of an astronaut riding a horse in photorealistic style). We asked the models to generate images with a target number of objects, in analogy with numerosity production tasks used in animal and human studies [14, 5]. The target objects belonged to the same classes used for the naming task, except for the “fast cards” category, which could be underrepresented in the corpora used to train these models. Examples of generated images are shown in the middle and bottom rows of Fig. 1.
Overall, DALL-E tended to generate images with white background and more definite objects, while Stable Diffusion adopted more artistic generation styles, producing either paintings, clip-arts, or realistic content. Similarly to the naming task, mean accuracy was low (DALL-E: 38.7%, Stable Diff: 33.3%), though DALL-E in some cases exhibited accurate response in the small number range (for the apples category, responses were error-free up to $n$ = 3). According to standard criteria, DALL-E would be considered a “Three”-knower, while Stable Diffusion would be considered a “Two”-knower [15]. Compared to the naming task, the response patterns were more homogeneous across categories (minimum correlation between CMs for DALL-E: 0.76, Stable Diff: 0.81).
Figure 2: Confusion matrices showing the distribution of models’ responses for different object categories. The x-axis represents the target number, while the y-axis represents the corresponding model responses. Response frequency is encoded using a grey scale (white = 0%, black = 100%). The first two rows refer to the numerosity naming task, which contains the additional “fast cards” category. The remaining rows refer to the numerosity production task, and contain two additional panels depicting the response distributions observed in behavioral studies with human adults [7] and children [16] across similar numerical ranges (note that for the adult data the maximum number of items to be estimated was eight). For the children data, the graph shows the response frequency of “Three”-knower children (larger circles indicate higher frequency), along with the estimated response likelihood (gray shading).
Adherence to Weber’s law
In humans and other animal species, numerosity estimation data collected through naming and production tasks approximately follow the law of scalar variability, as indicated by a constant coefficient of variation (standard error of the responses/mean response) [17, 18] (but see [19] for discussion of possible exceptions). To investigate whether models’ responses followed a scalar variability pattern, we used a well-established methodology implementing a log–log regression analysis [20]. If scalar variability holds, as the number $n$ increases the logs of the standard error $SE(n)$ and of the mean estimates $\mu(n)$ should form a straight line, with a slope $\beta=1$ [$log(SE(n))=\alpha+\beta log(\mu(n))$]. A slope that is significantly different from 1 would indicate that the relation between the standard errors and the means is a more complex power function, while a poor fit would more generally indicate that the response variability cannot be explained by linear or power trends.
We carried out a separate analysis for each model and for each object category. If a model exhibited error-free responses for small numbers, we excluded those numbers from the analysis. In almost half of the cases (8 out of 18), the regression analysis resulted in a poor fit ($p>0.05$), indicating that the distribution of responses did not systematically change as a function of numerosity, in sharp contrast with human and animal data. In the remaining cases the regression analysis was significant (all $p<0.05$; $r^{2}=0.64-0.96$), therefore we further tested whether the regression slopes were significantly different from 1. Results were mixed, and depended both on the model type and the object category.
In the numerosity naming task, the distribution of models’ responses did not follow the pattern observed in human estimation [7]. For ViLT, we observed scalar variability for the people (slope = 1.11, p > 0.05) and butterflies (slope = 1.14, p > 0.05) categories, while we observed power variability for the fast card category (slope = 1.37, p = 0.003). For BLIP2, the regression analysis never produced a good fit. In the numerosity production task the errors were more regularly distributed around the target number (also see Fig. 2). For DALL-E, we observed scalar variability for the dots category (slope = 0.82, p > 0.05), while we observed power variability for the apples (slope = 1.53, p = 0.024) and butterflies (slope = 1.49, p = 0.036) categories. For Stable Diffusion, we observed scalar variability for the apples (slope = 1.07, p > 0.05) and butterflies (slope = 1.15, p > 0.05) categories, while we observed power variability for the people (slope = 0.70, p = 0.030) and dots (slope = 0.60, p = 0.016) categories.
Overall, scalar variability was thus observed only in five specific conditions. In other five conditions the variability increased according to a power law, while in the remaining eight cases the response variability did not have a systematic trend.
Discussion
This work demonstrates that foundational AI models cannot reliably judge the number of objects in a visual scene, both in image-to-text and text-to-image tasks. Such a striking deficit is observed even for small sets, supporting a knower-level comparable to that of 3-year-old children [15, 16]. This might explain anecdotal evidence showing the failure of generative AI to synthesize realistic images featuring multiple instances of a body part, such as the correct number of fingers in a hand. Surprisingly, and in contrast with smaller-scale deep learning models [21, 22, 23], even approximate estimation was not robust and often failed to match the psychophysics of human numerosity perception. We believe that substantial progress in architecture design and training procedures is warranted to create AI systems that understand visual numerosity [24]. Grounding numeracy development in the visual number sense, as in the case of humans [3], might also be the key to enabling AI systems to acquire and fully master numerical and mathematical knowledge [25].
Materials and Methods
In the naming task, for each object class and target number we probed the models using 50 high-resolution ($1024\times 1024$ pixels) images created by randomly placing items of variable size on a uniform white background, with no overlap. Fast card stimuli were created using clip-arts of common objects (apples, bells, butterflies, candies, cars, fish, flowers, planes, stars) drawn in different colors (black, blue, green, orange, red). Responses without numerical content or with vague quantification (e.g., “a few”) were automatically discarded, but we ensured that at least 20 acceptable answers were produced for each class/number combination. In the production task, for each condition we probed the models to generate 50 high-resolution images, which were then manually labeled by one of the authors (K.H.) and individually double checked by another author. Images with an ambiguous number of objects or with ill-formed content were discarded (see Section Analysis of model responses for details). We ensured that at least 20 acceptable images were produced for each test condition.
Models
We considered two different Visual Question Answering (VQA) models. We used the vilt-b32-mlm version of the Vision-and-Language Transformer (ViLT) model [9] available through Hugging Face. This architecture incorporates text embeddings into a Vision Transformer (ViT), allowing it to have a minimal design for Vision-and-Language Pre-training and thus speeding-up model training and inference phases. It has a total of 87.4 million parameters. We also used the blip2-flan-t5-xl version of the Bootstrapping Language-Image Pre-training with Frozen Image Encoders and Large Language Models (BLIP-2) model [10], developed by Salesforce and available through Hugging Face. This architecture is considered an improved version of BLIP [26] and it is available in different versions, depending on the type and size of the backbone models. We explored all versions except for the t5-xxl model (due to GPU memory constraints) and found that the version using Flan-T5 as a language model yielded the best accuracy. The chosen model version has a total of 4.1 billion parameters.
We considered two different image generation models. We probed the DALL-E (version 2) model [11] using the API interface provided by OpenAI. DALL-E 2 is an improved version of the original text-to-image DALL-E model, featuring a total of 3.5 billion parameters. We also used the Stable Diffusion (version 2.1) model [12], developed by Stability AI and available through Hugging Face. Stable Diffusion v2.1 is a text-to-image latent diffusion model that combines an autoencoder with a diffusion model that is trained in the latent space of the autoencoder. It has approximately 500 million parameters.
Prompting methods
For each model, we explored different prompting methods and selected the one leading to the best performance, measured as mean absolute distance from the target number.
For the numerosity naming task, we implemented three different prompting methods. The first required to explicitly estimate the number of objects belonging to a specific category (i.e., How many apples / butterflies / people / dots / shapes are there in the picture?). The other two used the more general “objects” or “things” words to identify the items to count. For ViLT, the best performance was achieved with the generic “things” prompt, while for BLIP-2 the best performance was achieved with the category-specific prompts. To make sure that the models were prompted correctly, we carried out a control simulation related to a non-numerical task using the entire set of “apples” stimuli. We probed the VQA models with the following prompt: What does the image represent? and we considered as correct the following answers: apple(s) and fruit. All models almost always provided the correct answer across the entire set of stimuli (accuracy for ViLT : 96.8%, BLIP-2 : 100%), thus demonstrating a proper understanding of the image content and the prompt structure.
For the numerosity production task, both DALL-E and Stable Diffusion were initially prompted with the following text: An image with n apples / butterflies / people / dots (where n varied between 2 and 10). When n = 1 the prompt was adjusted to the singular form. However, for the dots category this prompting method resulted in poor generations: we obtained better results when the models were prompted with a more specific description of the image: An image with white background with n filled dots.
Analysis of model responses
For each model, we assessed the knower-level by applying standard criteria used in the literature on the development of counting skills [15] to the average responses across categories. To be considered an “$n$”-knower (i.e., “One”-knower, “Two”-knower, “Three”-knower, “Four”-knower) the model had to: 1) Give $n$ objects at least 67% of the time when asked for that number; and 2) Give $n$ objects no more than half as often when asked for a different number.
Responses given by VQA models were automatically parsed. If present, number words were converted to numerical values using the word2number Python library. The response was discarded if it contained multiple numbers or if it contained vague quantification terms (e.g., “a few”, “a bunch of”). We checked that at least 20 correct trials were recorded for each number/category. All models achieved the minimum number of tests required without the need to further prompt them (total number of responses discarded for ViLT: 0; BLIP-2: 6).
Images produced by the image generation models were discarded when more than 50% of at least one target object fell outside the margin, or the model generated only objects of non-target categories, or the image contained more than 20 objects. Target objects were not counted when they appeared at the border of the picture and only a very small portion (no more than 20%) of the object was visible, or they were not well-formed (e.g., the shape was clearly distorted) preventing unambiguous identification, or they were occluded by other objects, preventing unambiguous identification. Target objects were considered valid and thus counted when they appeared behind other objects but could still be reliably identified, or they were not entirely visible (e.g., more than 50% of the object mass fell outside the margin or behind other objects) but a salient feature of the object was clearly visible (e.g., the face of a person or the stalk of an apple), or they were represented in an unconventional way (e.g., an apple cut in half).
For DALL-E, the minimum number of trials was not achieved for the dots category with n = 10 (15), since the model in these cases most frequently generated more than 30 items. For Stable Diffusion, the minimum number was not achieved for the dots category with n = 1 (15), n = 8 (14), n = 9 (14) and n = 10 (13). Also in this case, for these prompts the model most frequently generated more than 30 items.
Software and computing hardware
All simulations and analyses were implemented using Python v3 and Google Colab. To test larger-scale models (BLIP-2 family) we used a virtual machine with an NVIDIA L4 GPU and 64 GB of RAM memory, which was allocated using the Google cloud computing platform. Data and code used in the current study are available on GitHub.
Acknowledgements
We are grateful to OpenAI for granting research access to the DALL-E APIs. This work was partially supported by the Italian Ministry of University and Research (PRIN grant n. C53D23004110006 to M.Z.).